Check whether or not the function f, defined on R by


f(x) = { 3x^2sin(1/2x), when x≠0


{ 0 ,when x=0


is derivable on R. If it is, is f' continuous at x=0? If f is not derivable , then define a derivable function on R

1
Expert's answer
2022-06-07T00:27:54-0400

Apply the Squeeze Theorem:


"-1\le\sin(1/2x)\le1, x\in \R"

Then


"-3x^2\le3x^2\sin(1/2x)\le3x^2, x\in \R"

We see that


"\lim\limits_{x\to0}(-3x^2)=0=\lim\limits_{x\to0}(3x^2)"

Then by the Squeeze Theorem


"\lim\limits_{x\to0}(3x^2\sin(1/2x))=0"

"\lim\limits_{x\to0}(3x^2\sin(1/2x))=0=f(0)"

The function "f(x)" is continuous on "\R."


Apply the Squeeze Theorem:


"-1\le\sin(1/2h)\le1, h\in \R"

Then


"-3h\le3h\sin(1/2h)\le3h, h\in \R"

We see that


"\lim\limits_{h\to0}(-3h)=0=\lim\limits_{h\to0}(3h)"

Then by the Squeeze Theorem


"\lim\limits_{h\to0}(\dfrac{f(0+h)-f(0)}{h})=\lim\limits_{h\to0}(\dfrac{3h^2\sin(1/2h)}{h})"

"=\lim\limits_{h\to0}(3h\sin(1/2h))=0=f'(0)"

The function "f(x)" is derivable on "\R."


"f'(x)=6x\sin(1/2x)+3x^2\cos(1/2x)(-\dfrac{1}{2x^2})"

"=6x\sin(1/2x)-\dfrac{3}{2}\cos(1/2x)"

"\lim\limits_{x\to0}(\cos(1/2x))=\text{does not exist}"

If "x_n=\dfrac{1}{4\pi n}" then "\cos(1/2x_n)=1, n\to \infin"


If "x_n=\dfrac{1}{\pi+2\pi n}" then "\cos(1/2x_n)=0, n\to \infin"


Therefore "\lim\limits_{x\to0}(f'(x))=\text{does not exist}."

The function "f'(x)" is not continuous on "\R."